Write the perimeter formula of a rectangle, and then substitute the following conditions into the formula: The perimeter of a rectangluar garden is 100 feet, and the lengh of the garden in 12 feet more than the width. use x to represent the width.
Answer by checkley77(9412)

2L+2W=PERIMETER
L=X+12
W=X
2(X+12)+2X=100
2X+24+2X=100
4X=100-24
4X=76
X=76/4
X=19 FT IS THE WIDTH.
19+12=31 FOR THE LENGTH.
PROOF:
2*31+2*19=100
62+38=100
100=100